OpenStudy (anonymous):

Simplify 3c^4 over 18c^12. will give metal to best answer. A. 1 over 6c^8 B. 6 over c^8 C. c^8 over 6 D. c^16 over 6

jigglypuff314 (jigglypuff314):

3/18 simplifies to ? c^4 / c^12 = 1/ c^(12-4) = 1/ c^8 then combine the two :)

OpenStudy (anonymous):

umm, so it is A?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

ok, thanks :) (again)

jigglypuff314 (jigglypuff314):

you're welcome (I have school now, see ya later!) :)

OpenStudy (anonymous):

ok, see ya